4. Use the trapezoidal rule with n=6 to estimate the following definite integral. Ś 1 dy 1+y5
Answer #1

3 4. dy - S f(yay 1+ ys fly)= 1+y5 Here, b=3, a=on=6 2 :h=3 = y = 0, 7 = 2 M = 1, I = 3 , 44=2, 4 = 5 , 4 = 3 x = f (yo) = [L